OpenStudy (anonymous):

If f(x)= 3x^2 - 8 and g(x)= 2x evvaluate f(x) divided by g(x) for x=2

OpenStudy (anonymous):

f(2)/g(2) = ?

OpenStudy (anonymous):

f(2) = ? g(2) = ?

OpenStudy (anonymous):

f(x)= 3x^2 - 8 if x=2 then f(2) = 3(2)^2 -8

OpenStudy (anonymous):

so the answer would be 1 because 4/4 is one

OpenStudy (anonymous):

yep ^_^

OpenStudy (anonymous):

can you help me please